y=-3x+5
5x-4y=-3
Show me your work also please!

Answer by mananth(16946)

3 x + 1 y = 5 .............1
5 x -4 y = -3 .............2
Eliminate y
multiply (1)by 4
Multiply (2) by 1
12 x 4 y = 20
5 x -4 y = -3
Add the two equations
17 x = 17
/ 17
x = 1
plug value of x in (1)
3 x + 1 y = 5
3 + 1 y = 5
1 y = 5-3
1 y = 2
y = 2